Etna volcano - eruption update
Watch Etna live
Watch Etna liveYou can watch various webcams and seismic signals of Etna live from our new tool: webcams.volcanodiscovery.com/Etna
Update Tue 22 Apr 11:09Occasional weak explosions continue at the New SE crater: two very small events were observed today at 03:45 and 04:05 (local time).
Update Fri 11 Apr 10:40
Activity at New SE crater has stopped again
Update Wed 09 Apr 09:18After a few days of bad weather that have prevented detailed observations, it seems that all is quiet again.
Tremor is slightly above normal background levels.
A new short pause or the end of activity?
We can only wait to see what Etna has in store.
Activity resumes at Etna's New SE crater
Update Wed 02 Apr 11:45After a few days of apparent calm, new activity resumes on Etna: weak strombolian activity occurs at the NSEC and renewed small lava effusion from a fissure vent at the eastern base of the cone.
Seismic tremor is slowly increasing
Update Sun 30 Mar 09:56Since the end of lava effusion 3 days ago, Etna has remained all quiet and the recent flow field is now slowly cooling down.
Activity at New SE crater ends
Update Thu 27 Mar 15:54During the night, the lava effusion and persistent mild explosive activity from the New SE crater ceased, after being nearly continuously active for over two months.
If this is a true end or only a short pause to the latest eruptive phase which on 22 January remains to be seen. The tremor fell back to low levels in correspondence.
Update Sat 22 Mar 13:29
A new effusive vent opened two days ago within the collapse scar alimented a lava flow but it seems no longer active today.
The tremor does not show significant variations.
Slowly increasing strombolian activity
Update Thu 20 Mar 17:59
Update Tue 18 Mar 18:31
Small ash emissions from Bocca Nuova
Update Mon 10 Mar 19:05
Update Sat 08 Mar 09:49
Continuing lava flows from NSEC
Update Sun 02 Mar 09:47No significant changes have occurred at the volcano. Lava continues to flow from the vents at the eastern flank of the New SE crater and produces overlapping flows, typically a few 100 m long, near the rim of and on the upper slope into the Valle del Bove. A small surge in lava effusion occurred during the past hours and can be seen at the end of the following time-lapse:
Update Tue 25 Feb 18:28
Update Sun 23 Feb 18:19
Weak intermittent strombolian activity continued at the summit vent. Tremor has risen a bit over the past 24 hours, but overall is low.
Update Sun 16 Feb 13:35Activity has remained essentially unchanged. Weak lava effusion from the fissure vents at the eastern flank of the New SE crater as well as sporadic strombolian explosion continue, accompanied by low tremor levels.
Another small effusive vent on the NE base of the cone had formed and been briefly active yesterday evening, feeding a short-lived small lava flow. This vent seems no longer active today.
Phreatic explosion at Bocca Nuova Crater
Update Sat 15 Feb 13:23
Weak strombolian activity continues at the sumit vent of the New SEC
The lava flow is still alimented but effusion rate has dropped and only the upper portion of the flow is currently active.
The seismic signal shows medium-low tremor levels.
Pyroclastic flow on eastern flank this morning
Update Tue 11 Feb 14:47
Most likely, explosive lava-snow interaction and destabilization of this area where new magma was pushing its way outside caused a sector to collapse and descend the Valle del Bove flank as a hot turbulent avalanche (=pyroclastic flow). It reached the bottom of Valle del Bove in less than 3 minutes and traveled about 3 km length, which implies an average speed of 60 km/h, with maximum speed probably in excess of 100 km/h (approx. 65 mph).
Here is the impressive video from INGV Catania:
See also Photos on facebook.
Update Tue 11 Feb 14:06The following is a comparison of some characteristic stages of the vents and lava flows on the eastern flank of the New SE cone:
Another new vent at the New SE crater
Update Tue 11 Feb 13:25Another new effusive vent opened up on the eastern side of the New SE crater around 04:00 this morning, erupting a new short-lived lava flow that stopped being active around 09:00.
The following time-lapse video shows the evolution of lava flows from the various vents, as well as pulses of strombolian activity since the start of the ongoing eruptive episode:
Update Mon 10 Feb 14:49
This morning at 10:20 (local time),after an ash explosion, a new vent opened over the vents that alimented the active lava flow.
Tremor is fluctuating, but there's a slow overall rising trend
Activity continues at the New SE crater
Update Sun 09 Feb 10:51Activity from the New SE crater continues essentially unchanged. Weak strombolian activity occurs at the summit vent, sometimes associated with more or less pronounced ash emissions.
When observed last evening, ejections at the NSEC rarely exceeded 50-100 m above the rim and only occasionally, bombs landed on the outer flank. The lava flow, only about 500-600 m long last night and half way down to the bottom of the Valle del Bove headwall, increased over night. It developed a second branch and its most advanced fronts have now reached again the base of the Valle del Bove.
Tremor fluctuates with an overall slowly rising trend. This seems to correspond rather to an increase in effusive(lava flow) than explosive activity.
Update Sat 08 Feb 09:40
Update Fri 07 Feb 09:16The lava flow and strombolian activity at the New SE crater have continued and activity has been increasing. The corresponding tremor signal is rising as well with a pronounced upwards trend over the past hours. It looks as if Etna is in for another paroxysm in the very near future.
Update Thu 06 Feb 11:54
The front of the old lava flow is stopped at 1660 m while the active lava flow is descending on the upper wall of the valle del Bove.
Ash emission continues from the summit vent of the New SE crater.
New lava flow and ash emissions from New SE crater
Update Tue 04 Feb 18:08After a short pause with likely no (or almost no) eruptive activity during 1-2 Feb, the volcano started renewed lava effusion from the fissure vent at the eastern foot of the New SE crater yesterday, followed by the onset of weak explosive activity(ash emissions) from the New SE crater this afternoon.
So far, this activity has been very weak and the lava flow progressed only a few 100 m. The tremor signal is low and doesn't show a clear trend at the moment.
Update Thu 30 Jan 12:42
Update Tue 28 Jan 09:57
Update Mon 27 Jan 01:35The fissure vent, at the eastern base of the New SE crater, has continued to produce active lava flows along the W wall of the Valle del Bove and, according to direct observations, its fronts have reached 1700 m elevation at NE of the "Monti Centenari" . In the afternoon a moderately dense pulsating ash column was emitted by weak, deep-seated strombolian activity at the summit vent.
The North-East Crater appeared to be covered by ash deposits and strong degassing activity continued from it.
Update Sun 26 Jan 10:18After a more vigorous phase yesterday, the effusion of the lava flow into the Valle del Bove and explosive (strombolian) activity at the New SE crater still continued this morning but have decreased in intensity.
Update Sat 25 Jan 14:04Last night the mild strombolian activity picked up to become intense.
The lava flow from the fissure vent at the eastern base of the New SE crater is still alimented and descend towards Valle del Bove.
After a steeply rising the volcanic tremor has dropped a bit but remains elevated.
Direct observations are difficult due to bad weather.
Update Fri 24 Jan 09:44The activity continued throughout the night with a constant lava effusion from the fissure vent at the eastern base of the New SE crater and with mild strombolian activity at the summit vent.
Seismic tremor continued to slowly but steadily increase.
Continuing lava flows
Update Thu 23 Jan 21:45Lava continues to flow from the fissure vent at the eastern base of the New SE crater towards the Valle del Leone. Mild strombolian explosions at the summit vent accompany this activity. The tremor and the intensity of this (still weak) activity are slowly increasing.
Increasing activity, lava flows from NSEC
Update Thu 23 Jan 09:27
The following time-lapse video from LAVE's webcam on Schiena dell'Asino (unfortunately out of focus during the night) shows the activity well.
At around 23:00 local time, 2 effusive vents on the fissure cutting through the eastern side of the cone open, feeding 2 small lava flows. Activity at the flank vents decreased (but not completely ceased) again during the early morning.
Strombolian activity at New SE crater
Update Wed 22 Jan 10:52
Update Thu 16 Jan 15:43
No ash emissions from the North-East Crater and the volcanic tremor signal is still low.
Update Sun 12 Jan 12:19
Update Sat 11 Jan 09:46Activity remains unchanged. The Northeast crater produces a weak ash plume currently drifting east. Tremor levels are fluctuating, but generally low.
Interestingly, a few shallow (6-7 km) earthquakes of magnitudes around 1.7-1.8 occurred along a W-E line over the past days.
Update Fri 10 Jan 11:49Activity and tremor have decreased for now, but weak ash emissions continue to be observed from the NE crater.
Update Thu 09 Jan 08:49The activity of the volcano has remained similar to the previous days: so far, relatively weak but more or less steady ash emissions are occurring at the North-East crater, and tremor continues its slow rise. Whether this is a precursor to more vigorous activity (from the New SE crater?) remains to be seen probably in the coming days.
Update Wed 08 Jan 12:22Weak sporadic ash emissions have been occurring from the NE crater since yesterday. Tremor is slightly elevated with an overall rising tendency.
Update Tue 06 Jan 20:18This morning, the North-East Crater appeared to be covered by ash deposits and strong, pulsating degassing activity continued from it. The volcanic tremor signal is still low.
Ash emissions from NE crater
Update Sat 04 Jan 11:24Again few days after the end of the last paroxysm at the New SE crater, weak ash emissions have started to occur from the North-East crater over night or this morning. Tremor is low.
Update Wed 01 Jan 13:13For the first day of the year, Etna is wearing a beautiful snow dress.
The tremor has descended back to low levels and lava effusion has almost completely ceased: only from the fissure on the lower NE flank of the New SE crater, a week effusive activity is continuing. Bocca Nuova, Voragine and the NE crater show intense degassing.
More on VolcanoDiscovery: